OpenStudy (anonymous):

A rumour spread exponentially through a college. 100 people have herad it by noon, and 200 people by 1 pm. How many people have heard it by: (a) 3pm (b) 12.30 pm (c) 1.45 pm?

OpenStudy (anonymous):

doubles every hour at noon 100 at 1 200 at 2 400 at 3 800

OpenStudy (anonymous):

How would you write that in exponential form?

OpenStudy (anonymous):

12:30 is a bit tricker. it is \[100\times 2^{\frac{1}{2}}\]

OpenStudy (anonymous):

if it doubles every hour with initial value 100 then "t" hours starting at noon, formula would be \[100\times 2^t\]

OpenStudy (anonymous):

so for three hours it is \[100\times 2^3=800\] and for half an hour it is \[100\times 2^{\frac{1}{2}}\]

OpenStudy (anonymous):

OK. That makes sense.. and the last would be 200 x 2^3/4? Or am I wrong.. just trying to get the ahng of it...

OpenStudy (anonymous):

yes you could write it like that

OpenStudy (anonymous):

Thanks :) I'm beginning to understand it now!

OpenStudy (anonymous):

You could write it as \( 100 \times 2^t \) where t is the elapsed time in *hour*.

OpenStudy (anonymous):

or you could write \[100\times 2^{\frac{7}{4}}\] which is the same thing.

OpenStudy (anonymous):

Ok, thanks again!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

It's 45 minutes passed 1:00.. or 15 minutes to 2:00.. so it's either ratio 7/4 or 3/4 the other way round

OpenStudy (anonymous):

Actually the answer should be 7/4 and not 3/4

OpenStudy (anonymous):

It gives the same answer :)

OpenStudy (anonymous):

it is the same thing

OpenStudy (anonymous):

Mea culpa, I thought you are multiplying by 100 both times.

OpenStudy (anonymous):

\(200 \times 2^\frac 34 = 100 \times 2 \times 2^\frac 34 = 100 \times 2^\frac 74 \)

OpenStudy (anonymous):

you can say "at 1 it was 200 and 3/4 of an hour later it is \[200\times 2^{\frac{3}{4}}\] or at noon it was 100 so 7/4 of an hour later it is \[100\times 2^{\frac{7}{4}}\] laws of exponents tell you they have to be equal because \[100\times 2^{\frac{7}{4}}=100\times 2^1\times 2^{\frac{3}{4}}=200\times 2^{\frac{3}{4}}\]

OpenStudy (anonymous):

what ffm said

OpenStudy (anonymous):

Yup, I kinda figured ;) Thanks anyway.